OpenStudy (anonymous):

Jill makes and sells 10 dozen chocolate chip cookies every day she sells them in bags of 5 for $2.00. How much more money will she make if she sells all of them in bags of 4 for $2.00?

OpenStudy (anonymous):

10 dozen 120 now bags of 5 means 120/5 = 24 bags 24 * 2.00$ = 48$ if, instead she will make bags of 4 120/4 = 30 bags 30*2.00$ = 60$ 12$ more

OpenStudy (toast):

10x12 = 120 120/5=24x2.00=48 120/4=30x2=60

OpenStudy (anonymous):

but the choice said A. $20. B.$12. C $6.00 D.$4.

OpenStudy (anonymous):

as i said 12$ more

OpenStudy (anonymous):

60 - 48

OpenStudy (toast):

I concur :) Answer is B) 12 dollars